# Machine learning meets $\mathfrak{su}(n)$ Lie algebra: Enhancing quantum dynamics learning with exact trace conservation

### v1

@@ Abstract
Machine learning (ML) has emerged as a promising tool for simulating quantum dissipative dynamics. However, existing methods often struggle to enforce key physical constraints, such as trace conservation, when modeling reduced density matrices (RDMs). While Physics-Informed Neural Networks (PINN) aim to address these challenges, they frequently fail to achieve full physical consistency. In this work, we introduce a novel approach that leverages the $\mathfrak{su}(n)$ Lie algebra to represent RDMs as a combination of an identity matrix and $n^{2}-1$ Hermitian, traceless, and orthogonal basis operators, where $n$ is the system’s dimension. By learning only the coefficients associated with this basis, our framework inherently ensures exact trace conservation, as the traceless nature of the basis restricts the trace contribution solely to the identity matrix. This eliminates the need for explicit trace-preserving penalty terms in the loss function, simplifying optimization and improving learning efficiency. We validate our approach on two benchmark quantum systems: the spin-boson model and the Fenna-Matthews-Olson complex. By comparing the performance of four neural network (NN) architectures—Purely Data-driven Physics-Uninformed Neural Networks (PUNN), $\mathfrak{su}(n)$ Lie algebra-based PUNN ($\mathfrak{su}(n)$-PUNN), traditional PINN, and $\mathfrak{su}(n)$ Lie algebra-based PINN ($\mathfrak{su}(n)$-PINN)—we highlight the limitations of conventional methods and demonstrate the superior accuracy, robustness, and efficiency of our approach in learning quantum dissipative dynamics.
@@ I Introduction
In recent years, machine learning (ML) has emerged as a powerful tool for learning complex spatio-temporal dynamics in high-dimensional spaces.[40, 41, 42, 43, 44, 40, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] It has demonstrated proficiency in predicting the future evolution of quantum states based on historical data, as well as directly forecasting quantum states as a function of time and/or simulation parameters.
@@ II.1 ML-enhanced quantum dissipative dynamics
II.2 Expansion of RDM in traceless basis
@@
III Results and Discussion
To showcase the effectiveness of our approach, we examine the relaxation dynamics of the SB model and exciton energy transfer (EET) in the FMO complex. We train three machine learning models: Purely Data-driven Physics-Uninformed Neural Networks (PUNN), $\mathfrak{su}(n)$-based PUNN ($\mathfrak{su}(n)$-PUNN), Physics-Informed Neural Networks (PINN), and $\mathfrak{su}(n)$-based PINN ($\mathfrak{su}(n)$-PINN). All models employ a hybrid architecture comprising convolutional neural networks (CNN), followed by LSTM layers, and fully connected dense layers (CNN-LSTM). Optimization was carried out using a composite loss function given, in general, by
@@ III Results and Discussion
and $\mathcal{L}_{5}$ constrains the eigenvalues to remain within the range $[0,1]$, a fundamental property of all realistic RDMs.
@@
Figure 1: Comparison of training and validation losses for PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN as a function of the number of epochs.
@@
To train the models, we employed OSTL-based dynamics propagation where RDM $\rho_{\rm{S}}(t)$ and coefficients $a_{i}$ at each time step was converted into a 1D vector. In both PUNN and PINN models, the Hermitian property $\mathbf{\rho}_{\rm{S},{ij}}(t)=\mathbf{\rho}_{\rm{S},{ji}}(t)^{*}$ is explicitly utilized. Consequently, only the diagonal elements (considering their real parts) and the upper off-diagonal elements were included in the analysis, with the real and imaginary components of the off-diagonal terms treated separately. In fact, because of this the loss-function term $\mathcal{L}_{3}$ is zero by definition and can be neglected. Further methodological details can be found in Ref.68. To optimize training process, farthest point sampling[69, 41] was applied to select a subset of trajectories. For each case of the SB model ($\varepsilon/\Delta=0$ and $1$) and the FMO complex (site-1 and 6), 400 trajectories were selected for training, and the remaining data were reserved for testing. Separate CNN-LSTM models were trained for the SB model and the FMO complex, each using an identical architecture. For comparison, we selected models with comparable training and validation losses.
@@
Figure 2: Comparison of trace conservation in quantum dissipative dynamics using PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN. The first column presents results for the asymmetric SB model, while the second column focuses on the FMO complex with initial excitation on site 1. For the SB model, the results correspond to an unseen dynamics characterized by $\varepsilon/\Delta=1.0$, $\gamma/\Delta=9.0$, $\lambda/\Delta=0.6$, and $\beta\Delta=1.0$. For the FMO complex, the parameters are $\gamma=400~{}\text{cm}^{-1}$, $\lambda=40~{}\text{cm}^{-1}$, and $T=90~{}\text{K}$. Details on training and prediction are provided in the Results and Discussion section.
@@
To evaluate the learning efficiency of all models, Fig. 1 presents the learning curves, illustrating the decrease in loss versus training epochs for each model. For the SB model, both PUNN and $\mathfrak{su}(n)$-PUNN exhibit comparable learning performance. However, $\mathfrak{su}(n)$-PINN demonstrates significantly superior learning efficiency across all epochs, with the loss function decreasing rapidly. In the case of the more complex FMO system, PUNN achieves a marginally lower loss compared to $\mathfrak{su}(n)$-PUNN. Nonetheless, similar to the SB model, $\mathfrak{su}(n)$-PINN outperforms PINN in terms of learning efficiency. These findings highlight that expanding the RDM in the $\mathfrak{su}(n)$ basis simplifies optimization and enhances learning efficiency and can describe more physically realistic behavior.
@@
Figure 3: Evaluation of positive semi-definiteness and eigenvalue constraints in the predicted RDMs across PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN models. Sparse positive eigenvalues are shown for clarity, with simulation parameters consistent with those in Fig. 3. In the SB model (left column), all models produce predominantly positive eigenvalues, though minor deviations are present in PUNN. For the FMO complex (right column), the purely data-driven $\mathfrak{su}(n)$-PUNN results in only 0.78% negative eigenvalues (11 out of 1407), outperforming PINN even with all constraints included. By incorporating the corresponding loss terms in $\mathfrak{su}(n)$-PINN, all predicted RDMs achieve complete positivity.
Figure 4: Evolution of the RDM’s elements, including population and coherence terms, as predicted by PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN models. The first column shows the relaxation dynamics for the SB model, while the second and third columns highlight the diagonal (population) and selected off-diagonal (coherence) elements of the RDM for the FMO complex, respectively. Simulation parameters are provided in the caption of Fig. 2 and accuracy comparison is given in Table 1.
Fig. 3 evaluates the positive semi-definiteness and eigenvalue constraints of the predicted RDMs by analyzing their eigenvalues. For simpler systems, such as the SB model, all models perform similarly, with nearly all eigenvalues remaining positive, except for a few negative eigenvalues in the PUNN results. In more complex cases, like the FMO complex, the PUNN model produces approximately 13% negative eigenvalues (183 out of 1407), while PINN yields 3.6% negative eigenvalues (52 out of 1407), demonstrating that incorporating the loss terms $\mathcal{L}_{4}$ and $\mathcal{L}_{5}$ improves the positivity of the predicted RDMs. In contrast, the purely data-driven $\mathfrak{su}(n)$-PUNN results in only 0.78% negative eigenvalues (11 out of 1407), outperforming PINN even with all constraints included. Finally, by adding the corresponding loss terms in $\mathfrak{su}(n)$-PINN, all predicted RDMs achieve complete positivity. This further confirms that our $\mathfrak{su}(n)$ approach enhances learning efficiency.
@@
$\mathfrak{su}(n)$-PUNN 0.0020 (0.0011, 0.0013) 0.0056 (0.0019 0.00077)
PINN 0.0027 (0.0028, 0.0018) 0.0092 (0.0021 0.00072)
$\mathfrak{su}(n)$-PINN 0.0017 (0.0014, 0.0014) 0.0037 (0.0014 0.00071)
To further assess the performance of the four models, Table 1 compares the accuracy of PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN in predicting the evolution of RDM elements (population and coherence terms) for the SB model and the FMO complex, as illustrated in Fig. 4. For the SB model, all models provide accurate predictions; however, $\mathfrak{su}(n)$-PUNN and $\mathfrak{su}(n)$-PINN achieve the lowest errors for both diagonal and off-diagonal elements, demonstrating superior learning capabilities compared to PUNN and PINN. Although PUNN and PINN also perform well, their errors are marginally higher.
@@
IV Concluding Remarks
@@
In this work, we addressed the limitations of existing ML approaches for modeling RDM, particularly their challenges in learning efficiency and enforcing physical constraints such as trace conservation. Traditional methods often directly model the RDM, relying on nonlinear loss functions to maintain physical constraints, which can hinder optimization and lead to unphysical predictions.
To overcome these limitations, we introduced a novel approach that leverages the $\mathfrak{su}(n)$ Lie algebra to expand the RDM in terms of an identity matrix and $n^{2}-1$ operators, where $n$ is the system’s dimension. These operators are Hermitian, traceless, orthogonal, and satisfy the quadratic Casimir operator condition. The traceless nature of the operators ensures that only the identity matrix contributes to the trace of the RDM, inherently guaranteeing exact trace conservation. Unlike methods such as PINN, which often struggle to enforce trace conservation during training, our approach eliminates the need for explicit penalty terms in the loss function, simplifying optimization and enhancing training stability. Additionally, we believe that the Lie-algebra approach provides a promising direction for generalizing training data to systems of varying dimensions.
In summary, our $\mathfrak{su}(n)$ Lie algebra-based framework not only ensures exact trace conservation but also simplifies the learning process, reduces computational complexity, and enhances the scalability of ML models for quantum dynamics. This approach represents a significant step forward in the accurate and efficient modeling of quantum systems using machine learning.
V Acknowledgments

### v2

@@ Abstract
Machine learning (ML) has emerged as a promising tool for simulating quantum dissipative dynamics. However, existing methods often struggle to enforce key physical constraints, such as trace conservation, when modeling reduced density matrices (RDMs). While Physics-Informed Neural Networks (PINN) aim to address these challenges, they frequently fail to achieve full physical consistency. In this work, we introduce a novel approach that leverages the $\mathfrak{su}(n)$ Lie algebra to represent RDMs as a combination of an identity matrix and $n^{2}-1$ Hermitian, traceless, and orthogonal basis operators, where $n$ is the system’s dimension. By learning only the coefficients associated with the operators, our framework inherently ensures exact trace conservation, as the traceless nature of the operators restricts the trace contribution solely to the identity matrix. This eliminates the need for explicit trace-preserving penalty terms in the loss function, simplifying optimization and improving learning efficiency. We validate our approach on two benchmark quantum systems: the spin-boson model and the Fenna-Matthews-Olson complex. By comparing the performance of four neural network (NN) architectures—Purely Data-driven Physics-Uninformed Neural Networks (PUNN), $\mathfrak{su}(n)$ Lie algebra-based PUNN ($\mathfrak{su}(n)$-PUNN), traditional PINN, and $\mathfrak{su}(n)$ Lie algebra-based PINN ($\mathfrak{su}(n)$-PINN)—we highlight the limitations of conventional methods and demonstrate the superior accuracy, robustness, and efficiency of our approach in learning quantum dissipative dynamics.
@@ I Introduction
In recent years, machine learning (ML) has emerged as a powerful tool for learning complex spatio-temporal dynamics in high-dimensional spaces.[40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62] It has demonstrated proficiency in predicting the future evolution of quantum states based on historical data, as well as directly forecasting quantum states as a function of time and/or simulation parameters.
@@ II.1 ML-enhanced quantum dissipative dynamics
II.2 Expansion of RDM using traceless operators
@@
III Loss function, data preparation and training
To showcase the effectiveness of our $\mathfrak{su}(n)$-bases approaches, we examine the relaxation dynamics of the SB model and exciton energy transfer (EET) in the FMO complex. For comparison, we utilize four models: Purely Data-driven Physics-Uninformed Neural Networks (PUNN), $\mathfrak{su}(n)$-based PUNN ($\mathfrak{su}(n)$-PUNN), Physics-Informed Neural Networks (PINN), and $\mathfrak{su}(n)$-based PINN ($\mathfrak{su}(n)$-PINN). All models employ a hybrid architecture comprising convolutional neural networks (CNN), followed by long short-term memory (LSTM) layers, and fully connected dense layers (CNN-LSTM). Optimization was carried out using a composite loss function given, in general, by
@@ III Loss function, data preparation and training
where, for a negative eigenvalue, the term $-\mu_{i}(t)$ becomes positive, resulting in $\mathbf{\text{max}}(0,-\mu_{i}(t))=-\mu_{i}(t)$, while it evaluates to 0 otherwise. The subsequent loss term, $\mathcal{L}_{5}$ constrains the eigenvalues to remain within the range $[0,1]$, a fundamental property of all realistic RDMs.
@@
Figure 1: Comparison of training and validation losses for PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN as a function of the number of epochs. Panels A and B compare PUNN with $\mathfrak{su}(n)$-PUNN for the SB (Panel A) and FMO (Panel B) models, respectively. Panels C and D compare PINN with $\mathfrak{su}(n)$-PINN for the SB (Panel C) and FMO (Panel D) models, respectively.
@@
To train the models, we employed OSTL-based dynamics propagation where RDM $\rho_{\rm{S}}(t)$ and coefficients $a_{i}$ at each time step was converted into a 1D vector. In both PUNN and PINN models, the Hermitian property $\mathbf{\rho}_{\rm{S},{ij}}(t)=\mathbf{\rho}_{\rm{S},{ji}}(t)^{*}$ is explicitly utilized. Consequently, only the diagonal elements (considering their real parts) and the upper off-diagonal elements were included in the analysis, with the real and imaginary components of the off-diagonal terms treated separately. As a result, the Hermiticity loss term $\mathcal{L}_{3}$ is zero by definition in these models and does not affect training. While not strictly necessary in these cases, we retained it to maintain consistency across model types and for pedagogical clarity. Further methodological details can be found in the Supporting Information. To optimize training process, farthest point sampling[69, 41] was applied to select a subset of trajectories. For each case of the SB model ($\varepsilon/\Delta=0$ and $1$) and the FMO complex (site-1 and 6), 400 trajectories were selected for training, and the remaining data were reserved for testing. Separate CNN-LSTM models were trained for the SB model and the FMO complex, each using an identical architecture. For comparison, we selected models with comparable training and validation losses.
@@
Figure 2: Comparison of trace conservation in quantum dissipative dynamics using PUNN (A, E), $\mathfrak{su}(n)$-PUNN (B, F), PINN (C, G), and $\mathfrak{su}(n)$-PINN (D, H). The first column (A-D) presents results for the asymmetric SB model, while the second column (E-H) focuses on the FMO complex with initial excitation on site 1. For the asymmetric SB model, the results correspond to an unseen dynamics characterized by $\varepsilon/\Delta=1.0$, $\gamma/\Delta=9.0$, $\lambda/\Delta=0.6$, and $\beta\Delta=1.0$. For the FMO complex, the parameters are $\gamma=400~{}\text{cm}^{-1}$, $\lambda=40~{}\text{cm}^{-1}$, and $T=90~{}\text{K}$. Details on training and prediction are provided in the Results and Discussion section.
IV Results and Discussion
In this section, we demonstrate the results including trace conservation, eigenvalue spectra, predicted dynamics and accuracy for all the models. This section presents a comprehensive evaluation of the proposed models. We analyze and compare the performance of PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN models in terms of learning efficiency, trace conservation, positive semi-definiteness, and dynamical accuracy for both the SB model and the FMO complex. Additionally, we investigate the influence of NN architectures and explore an alternative variant of PUNN, termed TC-PUNN, which enforces trace conservation by design.
IV.1 Comparison of PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN
@@
We begin by comparing the performance of the four primary model formulations: PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN. Figure 1 shows the learning curves, highlighting the decrease in training loss over epochs for each model. For the SB model, both PUNN and $\mathfrak{su}(n)$-PUNN exhibit comparable learning performance. However, $\mathfrak{su}(n)$-PINN demonstrates significantly superior learning efficiency across all epochs, with the loss function decreasing rapidly. In the case of the more complex FMO system, PUNN achieves a marginally lower loss compared to $\mathfrak{su}(n)$-PUNN. Nonetheless, similar to the SB model, $\mathfrak{su}(n)$-PINN outperforms PINN in terms of learning efficiency. These findings highlight that expanding the RDM in the $\mathfrak{su}(n)$ basis simplifies optimization and enhances learning efficiency and can describe more physically realistic behavior.
@@
Figure 3: Evaluation of positive semi-definiteness and eigenvalue spectra of the predicted RDMs using PUNN (A, E), $\mathfrak{su}(n)$-PUNN (B, F), PINN (C, G), and $\mathfrak{su}(n)$-PINN (D, H) for the SB model (A–D) and the FMO complex (E–H). Sparse positive eigenvalues are shown for clarity, and the number of negative eigenvalues is indicated for each model. Simulation parameters are consistent with those in Fig. 2.
To evaluates the performance Fig. 3 evaluates the positive semi-definiteness and eigenvalue constraints of the predicted RDMs by analyzing their eigenvalues. For simpler systems, such as the SB model, all models perform similarly, with nearly all eigenvalues remaining positive, except for a few negative eigenvalues in the PUNN results. In more complex cases, like the FMO complex, the PUNN model produces approximately 13% negative eigenvalues (183 out of 1407), while PINN yields 3.6% negative eigenvalues (52 out of 1407), demonstrating that incorporating the loss terms $\mathcal{L}_{4}$ and $\mathcal{L}_{5}$ improves the positivity of the predicted RDMs. In contrast, the purely data-driven $\mathfrak{su}(n)$-PUNN results in only 0.78% negative eigenvalues (11 out of 1407), outperforming PINN even with all constraints included. Finally, by adding the corresponding loss terms in $\mathfrak{su}(n)$-PINN, all predicted RDMs achieve complete positivity. This further confirms that our $\mathfrak{su}(n)$ approach enhances learning efficiency.
@@
$\mathfrak{su}(n)$-PUNN 0.0020 (0.0011, 0.0013) 0.0056 (0.0019, 0.00077)
PINN 0.0027 (0.0028, 0.0018) 0.0092 (0.0021, 0.00072)
$\mathfrak{su}(n)$-PINN 0.0017 (0.0014, 0.0014) 0.0037 (0.0014, 0.00071)
To further assess the performance of the four models, Table 1 compares the accuracy of PUNN, $\mathfrak{su}(n)$-PUNN, PINN, and $\mathfrak{su}(n)$-PINN in predicting the evolution of RDM elements (population and coherence terms) for the SB model and the FMO complex, as illustrated in Fig. S1 (Supporting Information). For the SB model, all models provide accurate predictions; however, $\mathfrak{su}(n)$-PUNN and $\mathfrak{su}(n)$-PINN achieve the lowest errors for both diagonal and off-diagonal elements, demonstrating superior learning capabilities compared to PUNN and PINN. Although PUNN and PINN also perform well, their errors are marginally higher.
@@
Figure 4: Comparison of PUNN (A–C) and TC-PUNN (D–F) models for the asymmetric SB model. Panel A and D show trace conservation; panels B and E show eigenvalue spectra of the predicted RDMs; and panels C and F depict the time evolution of representative RDM elements, including both populations and coherences. Reference dynamics is overlaid as dots. The simulation parameters are given in the caption of Fig. 2.
Figure 5: Comparison of PUNN (panels A–D) and TC-PUNN (panels E–H) models for the FMO complex with initial excitation on site-1. Panels A and E show trace conservation; panels B and F show eigenvalue spectra of the predicted RDMs; panels C and G depict the time evolution of diagonal RDM elements (populations); and panels D and H depict the time evolution of selected off-diagonal RDM elements (coherences). Reference dynamics is overlaid as dots. The simulation parameters are given in the caption of Fig. 2.
IV.2 PUNN and PINN with explicit trace conservation
As with the $\mathfrak{su}(n)$-based models, exact trace conservation can also be enforced in traditional PUNN and PINN architectures by construction. It can be achieved by excluding one of the diagonal elements (typically the $n$th) during training, and reconstructing it using the constraint $\rho_{nn,\text{S}}=1-\sum_{i=1}^{n-1}\rho_{ii,\text{S}}$. This approach guarantees that the trace of the RDM remains exactly unity at all time steps. We demonstrate this strategy only for PUNN—referred to as PUNN with exact trace conservation (TC-PUNN)—in both the SB model and FMO complex, with results shown in Fig. 4 and Fig. 5, respectively.
To evaluate the performance of TC-PUNN, we compare it against the standard PUNN model. While TC-PUNN ensures strict trace preservation by construction, it does not necessarily lead to improved overall predictive accuracy—especially for larger systems like the FMO complex, as summarized in Table 2. In the SB model, which involves only a single diagonal element, the use of exact trace conservation yields a notable improvement in accuracy for the diagonal elements. However, in larger systems such as the FMO complex, omitting a diagonal element from training and inferring it via the trace condition does not significantly improve performance. The benefit of exact trace conservation is diminished as the complexity of the system increases.
When comparing the eigenvalues spectra of the predicted RDMs, TC-PUNN shows a reduction in the number of negative eigenvalues for the SB model, indicating better physical consistency. However, this advantage does not extend to the FMO complex; in fact, the number of negative eigenvalues increases in this case (see Fig. 5F), suggesting that enforcing exact trace conservation may come at the cost of overall matrix consistency.
It is also important to note that while the final diagonal element is constrained to enforce trace conservation, its predicted value can diverge significantly from the true value, especially when the errors in the remaining $n-1$ diagonal elements accumulate. Thus, although TC-PUNN guarantees trace conservation, it does not always lead to better accuracy or physical fidelity across all systems.
Table 2: Comparison of PUNN and TC-PUNN models based on the average MAE for the diagonal (Diag) and off-diagonal (Off-diag) elements of the RDM predicted for the SB model and FMO complex. The off-diagonal errors are computed as the average MAE for both the real and imaginary components of the RDM elements.
Model SB Model FMO Complex
Diag Off-diag Diag Off-diag
(Real, Imag) (Real, Imag)
PUNN 0.0032 (0.0026, 0.0027) 0.0074 (0.0031, 0.0014)
TC-PUNN 0.0019 (0.0026, 0.0017) 0.0080 (0.0031, 0.0013)
Figure 6: Comparison of trace conservation in quantum dissipative dynamics for the asymmetric SB model using the PUNN approach with different neural network architectures: (A) MLP, (B) LSTM, (C) CNN, and (D) CNN-LSTM. All models are evaluated on an unseen test case with parameters given in the caption of Fig. 2.
Table 3: Comparison of different NN model architectures based on the average MAE for the diagonal (Diag) and off-diagonal (Off-diag) elements of the RDM predicted by the PUNN approach using MLP, LSTM, CNN, and CNN-LSTM models for the SB model. The off-diagonal errors are calculated as the average MAE for both the real and imaginary components of the RDM elements.
Model Diag Off-diag
(Real, Imag)
MLP 0.0030 (0.0019, 0.0021)
LSTM 0.0019 (0.0016, 0.0016)
CNN 0.0032 (0.0035, 0.0028)
CNN-LSTM 0.0032 (0.0026, 0.0027)
IV.3 Impact of NN architecture: MLP, LSTM, CNN, and CNN-LSTM
To investigate the influence of NN architecture on the modeling of quantum dissipative dynamics, we compare four common architectures—Multilayer perceptron (MLP), LSTM, CNN, and a hybrid CNN-LSTM—within the PUNN framework. This comparison is conducted using the SB model, evaluating each architecture in terms of trace conservation, dynamical accuracy, and positive semi-definiteness.
As shown in Fig. 6, the LSTM architecture, which is well-suited for temporal sequence modeling, demonstrates superior performance in preserving the trace, with comparative small fluctuations throughout the dynamics. CNN-LSTM follows closely, while CNN and MLP show larger deviations, with MLP performing the worst in this regard.
For the positive semi-definiteness of the predicted RDMs (Fig. S2, Supporting Information), LSTM and CNN-LSTM again outperform the other architectures, followed by CNN and MLP. When assessing the predictive accuracy of RDM dynamics, depicted in Fig. S3 (Supporting Information) and quantified by MAE in Table 3, the LSTM model achieves the lowest errors. Interestingly, MLP performs better in dynamical accuracy than CNN-LSTM and CNN, despite being inferior in enforcing physical constraints.
These results suggest that while architectures like LSTM are more effective in capturing both physical constraints and accurate dynamics, there can be a disconnect between physical fidelity (e.g., trace conservation, positivity) and pure prediction accuracy—highlighting a trade-off that must be considered when selecting an appropriate neural network structure for quantum dynamics modeling.
V Final remarks and future directions
@@
In this work, we addressed key limitations of conventional ML approaches for modeling RDMs, particularly their challenges in training efficiency and the enforcement of physical constraints such as trace conservation. Traditional methods often rely on directly modeling the full RDM and impose physical constraints via complex nonlinear loss terms, which can complicate optimization and lead to unphysical predictions.
To overcome these issues, we proposed a novel framework based on the $\mathfrak{su}(n)$ Lie algebra, which expresses the RDM as a linear combination of the identity operator and $n^{2}-1$ orthonormal, Hermitian, and traceless generators. The traceless property of the generators guarantees that only the identity component contributes to the trace, thereby ensuring exact trace conservation by construction. Unlike PINN or traditional models that require explicit penalty terms in the loss function to enforce constraints, our approach simplifies optimization and improves training stability. Furthermore, the algebraic formulation offers a natural path for generalizing ML models to quantum systems of varying dimensionality.
In addition to the core models presented in the main text, we also explored an alternative formulation, TC-PUNN, which enforces exact trace conservation in a conventional PUNN setting by reconstructing the final diagonal element of the RDM. While this variant succeeds in guaranteeing trace preservation, it does not consistently improve predictive accuracy or physical consistency across all systems—particularly in higher-dimensional settings, where error propagation can degrade the overall fidelity of the predictions.
Moreover, we conducted a comparative study of different neural network architectures—MLP, LSTM, CNN, and CNN-LSTM—within the PUNN framework to assess the impact of network structure on performance. The results reveal that architectures with temporal modeling capabilities, particularly LSTM, generally outperform others in terms of trace conservation, eigenvalue positivity, and prediction accuracy. However, we also observed that better physical constraint satisfaction does not always correlate with higher accuracy in dynamics prediction, highlighting important trade-offs when designing neural networks for quantum modeling.
Looking ahead, several promising directions can be pursued to further advance this approach. One such avenue is exploring the ability to learn from smaller subsystems and extrapolate to larger ones—for example, training on $N<<M$ states and generalizing to $M>N$. Another valuable extension involves incorporating more complex and structured spectral densities to better capture realistic system-bath interactions. From an optimization perspective, replacing manually tuned loss weights with Lagrangian multipliers could offer a more principled and interpretable method for enforcing physical constraints, particularly within PINN-based architectures. Finally, the $\mathfrak{su}(n)$-based PINN framework employed in this work may also provide a natural foundation for tackling more advanced constraints, such as enforcing $n$-body positivity in systems with permutation symmetry—an important and currently underexplored area in quantum physics.
In summary, our Lie algebra-based ML framework offers a principled and scalable solution for modeling open quantum system dynamics. It inherently satisfies trace conservation, reduces reliance on complex loss terms, and facilitates more stable and efficient training. We believe this approach represents a significant step toward more accurate, generalizable, and physically consistent ML models for quantum dynamics.
VI Acknowledgments